# AirIO: Learning Inertial Odometry with Enhanced IMU Feature Observability air-io.github.io

### v1

@@ III-C AirIO Motion Network & Training
Based on the previous analysis, we introduce a separate encoding of the drone’s attitude to the network beyond the IMU encoder with body-frame representation, shown in Fig. 2(c). Inspired by [29], we map the drone’s orientation to the ${\mathfrak{so}(3)}$ Lie algebra space for compact, unconstrained, and continuous representation of 3D rotations and smoother network gradient to mitigate numerical instability.
We use CNN encoders to extract the features from the body-frame IMU data $(\mathbf{w}^{\mathcal{B}},\mathbf{a}^{\mathcal{B}})$ and the orientation of the drone $\boldsymbol{\xi}\in{\mathfrak{so}(3)}$. After concatenating the features from the attitude encoder and the IMU encoder, we employ bi-directional GRU layers to extract the latent feature $\mathbf{h}_{i}$ and model the temporal dependencies in the drone dynamics. In the decoder, we use two linear layers to output the body-frame velocity and its uncertainty. Overall the motion network can be written as: $(\hat{\mathbf{v}}^{\mathcal{B}},\hat{\boldsymbol{\sigma}}^{v})=f_{\boldsymbol{\theta}}(\mathbf{w}^{\mathcal{B}},\mathbf{a}^{\mathcal{B}},\boldsymbol{\xi})$. The model is supervised using an uncertainty loss inspired by [8]:
$L=(\mathbf{v}^{\mathcal{B}}_{i}-\hat{\mathbf{v}}^{\mathcal{B}}_{i})\hat{\boldsymbol{\Sigma}}^{v}_{i}{}^{-1}{(\mathbf{v}^{\mathcal{B}}_{i}-\hat{\mathbf{v}}^{\mathcal{B}}_{i}})^{T}+\ln(\det\hat{\boldsymbol{\Sigma}}^{v}_{i}),$ (3)
where $\hat{\boldsymbol{\Sigma}}^{v}_{i}$ are $3\times 3$ covariance matrix for the $i^{th}$ data. Similar to the setting in [8], we simplified the covariance as $\hat{\boldsymbol{\Sigma}}^{v}_{i}=\text{diag}(\hat{\boldsymbol{\sigma}}^{v}_{i}{}^{2})$, where $\hat{\boldsymbol{\sigma}}^{v}_{i}$ is learned uncertainty from the network. For orientation input $\boldsymbol{\xi}$, we use the ground truth orientations during training and replace them with EKF-estimated orientations during testing.
@@ III-D Extended Kalman Filter
Figure 6: AirIO system pipeline. A tightly coupled EKF that fuses the AirIO motion network with learning-based IMU preintegration network AirIMU. Instead of using constant uncertainty parameters, we employ learned uncertainty $\hat{\boldsymbol{\sigma}}^{v}$ from the AirIO network and the learned gyroscope uncertainty $\hat{\boldsymbol{\sigma}}^{g}$ and accelerometer $\hat{\boldsymbol{\sigma}}^{a}$ in our EKF system.
Different from the existing IO methods [8, 6] that employ a motion network to capture the relative translation. In our implementation, our network models the body-frame velocity. This setup disentangles the current state of the drone from the previous state, which simplifies the EKF structure of our model. The full state of the filter is $\mathbf{X}=(\mathbf{R}^{\mathcal{G}}_{i},\mathbf{v}^{\mathcal{G}}_{i},\mathbf{p}^{\mathcal{G}}_{i},\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i}})$, where $\mathbf{b}_{a_{i}}$ and $\mathbf{b}_{g_{i}}$ is the bias of the accelerometer and the gyroscope respectively. Following the common setup of the error-based filtering method, we define the error-state representation of the current filter state as $\delta\mathbf{X}=(\delta{\boldsymbol{\xi}}^{\mathcal{G}}_{i},\delta\mathbf{v}^{\mathcal{G}}_{i},\delta\mathbf{p}^{\mathcal{G}}_{i},\delta\mathbf{b}_{a_{i}},\delta\mathbf{b}_{g_{i}})$. For the rotation error, we define $\delta{\boldsymbol{\xi}}^{\mathcal{G}}_{i}=\mathbf{Log}(\mathbf{R}_{i}\hat{\mathbf{R}}_{i}^{-1})\in{\mathfrak{so}(3)}$, where $\mathbf{Log}(\cdot)$ denotes logarithm map operator.
AirIMU feature correction and uncertainty estimation To filter the noise and, more importantly, quantify the uncertainty of the IMU sensor model, we leverage AirIMU [29] to filter the raw IMU data. The model was trained through the differentiable IMU integrator and the covariance propagation to predict the corrections of the gyroscope and the accelerometer $[\boldsymbol{\sigma}_{g_{i}},\boldsymbol{\sigma}_{a_{i}}]\in\mathbb{R}^{3}$ with the corresponding uncertainty $[\boldsymbol{\eta}_{g_{i}},\boldsymbol{\eta}_{a_{i}}]\in\mathbb{R}^{3}$:
@@
We later filtered the IMU measurement by the predicted correction $\hat{\mathbf{a}}=\mathbf{a}+\hat{\boldsymbol{\sigma}}_{a_{i}},\hat{\mathbf{w}}=\mathbf{w}+\hat{\boldsymbol{\sigma}}_{g_{i}}$.
Filter Propagation The kinematic motion model of the filter propagation is:
$\begin{split}&\mathbf{R}_{i+1}=\mathbf{R}_{i}\cdot\mathbf{Exp}(\hat{\mathbf{w}}_{i}-\mathbf{b}_{g_{i}})\Delta t,\\ &\mathbf{v}_{i+1}=\mathbf{v}_{i}+\mathbf{R}_{i}(\hat{\mathbf{a}}_{i}-\mathbf{b}_{a_{i}}){\Delta}t,\\ &\mathbf{p}_{i+1}=\mathbf{p}_{i}+\mathbf{v}_{i}\Delta t+\frac{1}{2}\Delta t^{2}(\hat{\mathbf{a}}_{i}),\\ &\mathbf{b}_{a_{i+1}}=\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i+1}}=\mathbf{b}_{g_{i}},\end{split}$ (5)
Here, the $\mathbf{Exp}(\cdot)$ denotes mapping from the log space to exponential space. The linearized propagation model is:
@@
where the $\mathbf{W}_{i}$ is the covariance matrix of the $i$-th state, including the learned uncertainty of the IMU $\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}}$ and sensor bias random walk $\boldsymbol{\eta}_{b_{g}},\boldsymbol{\eta}_{b_{a}}$.
@@
$h(\mathbf{X})=\mathbf{R}_{i}^{T}\cdot\mathbf{v}_{i}=\hat{\mathbf{v}}_{i}+\hat{\boldsymbol{\eta}}_{v_{i}}.$ (8)
@@
$\begin{split}\mathbf{H}_{\delta v_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\mathbf{v}_{i}}=\mathbf{R}_{i}^{T}\\ \mathbf{H}_{\delta\xi_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\boldsymbol{\xi}_{i}}=\mathbf{R}_{i}^{T}[\mathbf{v}_{i}]_{\times}\end{split}$ (9)
Finally, the $\mathbf{H}$ and learned covariance are used in calculating the Kalman Gain:
$\begin{split}&\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{HPH}^{T}+\hat{\boldsymbol{\Sigma}}^{v}_{i})\\ &\mathbf{X}\leftarrow\mathbf{X}\oplus(\mathbf{K}(h(\mathbf{X})-\hat{\mathbf{v}}_{i}))\\ &\mathbf{P}\leftarrow(\mathbf{I}-\mathbf{KH})\mathbf{P}(\mathbf{I}-\mathbf{KH})^{T}+\mathbf{K}\hat{\boldsymbol{\Sigma}}^{v}_{i}\mathbf{K}^{T}\end{split}$ (10)
Operator $\oplus$ denotes the additional operation except for the orientation, where the update performs $\mathbf{R}\leftarrow\mathbf{Exp}(\boldsymbol{\xi})\cdot\mathbf{R}$.
@@ IV Experiment
In this paper, we conduct evaluations on three quadrotor datasets: the real-world EuRoC dataset [32], the challenging drone racing dataset BlackBird [30], and our custom simulation dataset collected using Pegasus Simulator [7]. The Blackbird dataset is an aggressive indoor flight dataset. Compared to the EuRoC dataset, Blackbird presents more aggressive maneuvers and high-speed dynamics, making estimating the drone’s motion more challenging. The Pegasus dataset enables ablation studies with eliminating the ground-truth inaccuracies in real-world settings.
We compare AirIO with several methods. First, we select model-based IO methods like Baseline [14], the IMU preintegration approach in raw IMU data, and AirIMU [29], a hybrid method that corrects the raw IMU noise with data-driven method and integrates it with the IMU kinematic function. For learning-based IO, we compare against end-to-end trained model RoNIN [9] and an EKF-fused algorithm TLIO [8]. For recent learning-based IO methods targeting to UAVs, we compare against IMO [6], a control signal embedded IO specifically designed for drone racing. It incorporates additional thrust signals to capture drone motion.
IV-A Metrics Definition
We use the following metrics for evaluation:
Absolute Translation Error (ATE, $\mathrm{m}$)
calculates the average Root Mean Squared Error (RMSE) between the estimated and ground-truth positions over all time points.
Relative Translation Error (RTE, $\mathrm{m}$)
calculates the average RMSE of the relative displacements over predefined time intervals. Our evaluation specifically adopts a 5-second interval configuration.
IV-B Training Details
We use the Adam optimizer with an initialized learning rate of 0.001. The learning rate scheduler is implemented following the ReduceLROnPlateau, with patience of 5 epochs and a decay factor of 0.2. The batch size is 128. Our training and testing window size is set to 1000 frames (corresponding to 5 seconds in the EuRoC dataset). In the CNN encoder, we include a dropout layer with $p=0.5$ to reduce overfitting.
@@ IV-B Training Details
Seen
    Clover 15.769 2.027 21.696 1.701 3.074 1.367 1.464 0.797 0.381 0.681 0.553 0.398 0.599 0.642
    Egg 66.170 7.677 13.675 4.178 2.449 2.552 2.227 2.398 1.153 0.828 0.649 0.310 0.549 0.307
    halfMoon 19.165 3.386 15.822 1.524 1.263 0.864 0.956 0.475 0.761 0.240 0.378 0.304 0.535 0.327
    Star 20.490 4.556 13.099 4.782 3.150 3.266 0.680 0.784 2.130 3.066 0.672 0.373 0.591 0.468
    Winter 15.781 2.395 11.800 1.772 1.031 1.089 0.616 0.755 0.219 0.206 0.396 0.183 0.332 0.196
Avg. 27.475 4.008 15.218 2.791 2.193 1.828 1.189 1.042 0.929 1.004 0.530 0.314 0.521 0.388
Unseen
    Ampersand 41.138 5.026 24.060 4.859 5.467 5.321 4.665 4.078 17.664 10.039 2.375 1.148 2.346 1.212
    Sid 13.108 2.055 16.601 2.782 18.581 10.665 7.323 8.427 9.967 6.992 0.683 0.539 0.686 0.531
    Oval 15.365 2.400 13.367 2.320 20.399 12.12 1.276 1.045 5.670 2.863 0.987 0.569 0.768 0.572
    Sphinx 3.429 2.713 10.737 3.736 11.537 7.762 2.005 2.408 5.629 5.395 1.204 1.020 1.119 1.017
    BentDice 28.307 2.342 38.060 3.036 11.453 6.792 2.119 1.747 6.143 4.519 1.156 0.955 1.111 1.226
Avg. 20.269 2.907 20.565 3.347 13.487 8.532 3.478 3.541 9.015 5.962 1.281 0.846 1.206 0.912
Dead-reckoning IMU pre-integration.
IV-C System Performance
Blackbird Dataset Following IMO [6], we employ the same sequences from Blackbird and adopt their same train-test split configuration. We term these sequences as SEEN sequences because the training and testing data are from the same sequence. To evaluate generalization, we select five additional trajectories from BlackBird dataset as UNSEEN sequences, which are excluded from the training set. You may find more details on separation in Appendix. A-A1.

### v2

@@ III-C AirIO Motion Network & Training
Based on the previous analysis, we introduce a separate encoding of the drone’s attitude $\boldsymbol{\xi}\in{\mathfrak{so}(3)}$ to the motion network beyond the IMU encoder with body-frame representation, shown in Fig. 1(c). After fusing this information, the network predicts the body-frame velocity ${}^{\mathcal{B}}\hat{\mathbf{v}}$ and the corresponding uncertainty $\hat{\boldsymbol{\eta}}^{v}$. The motion network can be written as: $({}^{\mathcal{B}}\hat{\mathbf{v}},\hat{\boldsymbol{\eta}}^{v})=f_{\boldsymbol{\theta}}({}^{\mathcal{B}}\mathbf{w},{}^{\mathcal{B}}\mathbf{a},\boldsymbol{\xi})$.
Inspired by [30], we map the drone’s orientation to the ${\mathfrak{so}(3)}$ Lie algebra space for compact and continuous representation of 3D rotations and smoother network gradient to mitigate numerical instability. We then leverage a CNN encoder to encode ${\mathfrak{so}(3)}$.
After concatenating the features from the attitude encoder and the IMU encoder, we employ bi-directional GRU layers to extract the latent feature $\mathbf{h}_{i}$ and model the temporal dependencies in the drone dynamics. We leverage this feature to estimate the body-frame velocity and the corresponding uncertainty by two MLP decoders.
To jointly supervise the velocity and the corresponding uncertainty ${}^{\mathcal{B}}\hat{\mathbf{v}}_{i}$, we designed a combined loss function:
$\mathcal{L}=\mathcal{L}_{\text{Huber}}+\lambda\mathcal{L}_{C},$ (3)
where the $\lambda=1e^{-4}$. The $\mathcal{L}_{\text{Huber}}$ is the Huber loss between the predicted body-frame velocity ${}^{\mathcal{B}}\mathbf{v}_{i}$ and the body-frame ground-truth velocity ${}^{\mathcal{B}}\hat{\mathbf{v}}_{i}$, defined as:
$\mathcal{L}_{\text{Huber}}=\begin{cases}\frac{1}{2}({}^{\mathcal{B}}\mathbf{v}_{i}-{}^{\mathcal{B}}\hat{\mathbf{v}}_{i})^{2}&\text{if }|{}^{\mathcal{B}}\mathbf{v}_{i}-{}^{\mathcal{B}}\hat{\mathbf{v}}_{i}|<\delta\\ \delta\cdot\left(|{}^{\mathcal{B}}\mathbf{v}_{i}-{}^{\mathcal{B}}\hat{\mathbf{v}}_{i}|-\frac{1}{2}\delta\right)&\text{otherwise}\end{cases},$ (4)
where we set the $\delta=0.005$.
To supervise the uncertainty of the velocity, we employ an uncertainty loss inspired by [7], assuming that the estimated velocity uncertainty follows a Gaussian distribution:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{C}}=({}^{\mathcal{B}}\mathbf{v}_{i}-{}^{\mathcal{B}}\hat{\mathbf{v}}_{i})\hat{\boldsymbol{\Sigma}}^{v}_{i}{}^{-1}{({}^{\mathcal{B}}\mathbf{v}_{i}-{}^{\mathcal{B}}\hat{\mathbf{v}}_{i}})^{T}+\ln(\det\hat{\boldsymbol{\Sigma}}^{v}_{i}),$ (5)
where $\hat{\boldsymbol{\Sigma}}^{v}_{i}$ is $3\times 3$ covariance matrix for the $i^{th}$ data. Similar to the setting in [7], we simplified the covariance as $\hat{\boldsymbol{\Sigma}}^{v}_{i}=\text{diag}(\hat{\boldsymbol{\eta}}^{v}_{i}{}^{2})$, where $\hat{\boldsymbol{\eta}}^{v}_{i}$ is learned uncertainty from network.
For orientation $\boldsymbol{\xi}$, we use the ground truth during training and replace it with EKF-estimated values during testing.
@@ III-D Extended Kalman Filter
Figure 5: AirIO system pipeline. A tightly coupled EKF that fuses the AirIO motion network with a learning-based IMU preintegration network AirIMU. Instead of using constant uncertainty parameters, we employ learned uncertainty $\hat{\boldsymbol{\eta}}^{v}$ from the AirIO network and the learned gyroscope uncertainty $\hat{\boldsymbol{\eta}}^{g}$ and accelerometer $\hat{\boldsymbol{\eta}}^{a}$ from the AirIMU in our EKF system.
Different from the existing IO methods [7, 6] which employ a motion network to capture relative translation and maintain historic states to optimize relative pose for state pairs. AirIO network models the body-frame velocity. As a result, we only need to constrain the velocity of the current frame. which significantly reduces the number of state parameters and alleviates the need for frame culling. This setup isolates the current state of the drone from the previous state, which simplifies the EKF structure. The full state of the filter is $\mathbf{X}=(\mathbf{R}_{i},{}^{\mathcal{G}}\mathbf{v}_{i},{}^{\mathcal{G}}\mathbf{p}_{i},\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i}})$, where $\mathbf{b}_{a_{i}}$ and $\mathbf{b}_{g_{i}}$ are the bias of the accelerometer and the gyroscope respectively. Following the common setup of the error-based filtering method, we define the error-state of the current filter as $\delta\mathbf{X}=(\delta{{}\boldsymbol{\xi}}_{i},\delta{}^{\mathcal{G}}\mathbf{v}_{i},\delta{}^{\mathcal{G}}\mathbf{p}_{i},\delta\mathbf{b}_{a_{i}},\delta\mathbf{b}_{g_{i}})$. For the rotation error, we define $\delta{\boldsymbol{\xi}}_{i}={\log}_{SO3}(\mathbf{R}_{i}\hat{\mathbf{R}}_{i}^{-1})\in{\mathfrak{so}(3)}$, where ${\log}_{SO3}(\cdot)$ denotes logarithm map operator.
AirIMU feature correction and uncertainty estimation To filter the noise and, more importantly, quantify the uncertainty of the IMU pre-integration, we leverage AirIMU [30] to pre-process the raw IMU data. This model estimates the corrections of the gyroscope and the accelerometer, and also the uncertainty of the IMU sensor:
@@
where the $\mathbf{\theta}$ is the parameter of the neural network. The AirIMU network is trained by a differentiable IMU integrator and covariance propagator. We later filtered the IMU measurement by the predicted correction $\hat{\mathbf{a}}=\mathbf{a}+\hat{\boldsymbol{\sigma}}_{a_{i}},\hat{\mathbf{w}}=\mathbf{w}+\hat{\boldsymbol{\sigma}}_{g_{i}}$. In the EKF system, the learned uncertainty $(\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}})$ is leveraged in the sensor covariance modeling.
Filter Propagation The kinematic motion model is:
$\begin{split}&\mathbf{R}_{i+1}=\mathbf{R}_{i}\cdot\text{Exp}(\hat{\mathbf{w}}_{i}-\mathbf{b}_{g_{i}})\Delta t,\\ &\mathbf{v}_{i+1}=\mathbf{v}_{i}+\mathbf{R}_{i}(\hat{\mathbf{a}}_{i}-\mathbf{b}_{a_{i}}){\Delta}t,\\ &\mathbf{p}_{i+1}=\mathbf{p}_{i}+\mathbf{v}_{i}\Delta t+\frac{1}{2}\Delta t^{2}\mathbf{R}_{i}(\hat{\mathbf{a}}_{i}-\mathbf{b}_{a_{i}}),\\ &\mathbf{b}_{a_{i+1}}=\mathbf{b}_{a_{i}},\mathbf{b}_{g_{i+1}}=\mathbf{b}_{g_{i}},\end{split}$ (7)
Here, the $\text{Exp}(\cdot)$ denotes mapping from the log space to exponential space. The linearized propagation model is:
@@
where the $\mathbf{W}_{i}$ is the covariance matrix of the $i$-th input, including the learned uncertainty of the IMU $\hat{\boldsymbol{\eta}}_{g_{i}},\hat{\boldsymbol{\eta}}_{a_{i}}$ and sensor bias random walk $\boldsymbol{\eta}_{b_{g}},\boldsymbol{\eta}_{b_{a}}$.
@@
$h(\mathbf{X})=\hat{\mathbf{R}}_{i}^{T}\cdot\mathbf{v}_{i}=\hat{\mathbf{v}}_{i}+\hat{\boldsymbol{\eta}}_{v_{i}}.$ (10)
@@
$\begin{split}\mathbf{H}_{\delta v_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\mathbf{v}_{i}}=\hat{\mathbf{R}}_{i}^{T}\\ \mathbf{H}_{\delta\xi_{i}}&=\frac{\partial h(\mathbf{X})}{\partial\delta\boldsymbol{\xi}_{i}}=\hat{\mathbf{R}}_{i}^{T}[\mathbf{v}_{i}]_{\times}\end{split}$ (11)
Finally, the Kalman Gain and the update can be computed:
$\begin{split}&\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}HPH}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\boldsymbol{\Sigma}}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}i}})^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-1}}\\ &\mathbf{X}\leftarrow\mathbf{X}\oplus(\mathbf{K}(h(\mathbf{X})-\hat{\mathbf{v}}_{i}))\\ &\mathbf{P}\leftarrow(\mathbf{I}-\mathbf{KH})\mathbf{P}(\mathbf{I}-\mathbf{KH})^{T}+\mathbf{K}\hat{\boldsymbol{\Sigma}}^{v}_{i}\mathbf{K}^{T}\end{split}$ (12)
Operator $\oplus$ denotes the additional operation except for the orientation, where the update performs $\mathbf{R}\leftarrow\text{Exp}(\boldsymbol{\xi})\cdot\mathbf{R}$.
@@ IV Experiment
Seen Clover 15.769 2.027 6.163 0.294 3.074 1.367 1.464 0.797 0.381 0.681 0.434 0.368 0.367 0.391
Egg 66.17 7.677 13.293 3.801 2.449 2.552 2.227 2.398 1.153 0.828 0.713 0.391 0.408 0.344
halfMoon 19.165 3.386 4.174 0.746 1.263 0.864 0.956 0.475 0.761 0.24 0.491 0.256 0.457 0.253
Star 20.49 4.556 4.347 1.933 3.15 3.266 0.68 0.784 2.13 3.066 0.442 0.401 0.477 0.341
Winter 15.781 2.395 7.445 0.743 1.031 1.089 0.616 0.755 0.219 0.206 0.348 0.147 0.307 0.164
Avg. 27.475 4.008 7.084 1.503 2.193 1.828 1.189 1.042 0.929 1.004 0.486 0.312 0.403 0.299
Unseen Ampersand 41.138 5.026 24.738 4.379 5.467 5.321 4.665 4.078 17.664 10.039 2.303 1.145 2.334 1.154
Sid 13.108 2.055 16.832 2.131 18.581 10.665 7.323 8.427 9.967 6.992 0.717 0.527 0.874 0.49
Oval 15.365 2.4 8.312 1.553 20.399 12.12 1.276 1.045 5.67 2.863 0.976 0.583 0.828 0.54
Sphinx 3.429 2.713 2.828 2.024 11.537 7.762 2.005 2.408 5.629 5.395 1.145 1.008 1.178 1.033
BentDice 28.307 2.342 54.261 4.39 11.453 6.792 2.119 1.747 6.143 4.519 1.360 1.000 1.331 0.955
Avg. 20.269 2.907 21.403 2.896 13.487 8.532 3.478 3.541 9.015 5.962 1.300 0.853 1.309 0.834
Leverage thrust information as the input data for training and inference.
Leverage ground truth orientation to transform the input data. Dead-reckoning IMU pre-integration.
IV-A Experiment Setup
Datasets We evaluate the proposed approach on two quadrotor datasets: the real-world EuRoC dataset [31] and the challenging drone racing dataset Blackbird [27]. EuRoC datasets are a widely used benchmark for evaluating odometry. Blackbird dataset is an indoor flight dataset, presenting more aggressive maneuvers and high-speed dynamics. For ablation studies, we additionally include our custom simulated dataset collected using the NVIDIA IsaacSim simulator with Pegasus autopilot [28]. The Pegasus dataset also presents diverse maneuvers but in the ideal environment without external disturbance, allowing us to work with cleaner data for ground-truth accuracy.
Baseline Several state-of-the-art methods are selected for comparison. For model-based IO, we include Baseline [13], the IMU preintegration approach in raw IMU data, and AirIMU [30], a hybrid method that corrects the raw IMU noise with a data-driven method and integrates it with the IMU kinematic function. For learning-based IO, we compare against an end-to-end trained model RoNIN [8] and an EKF-fused algorithm TLIO [7]. For learning-based IO methods designed for UAVs, we evaluate IMO [6], a control signal-embedded IO specifically developed for drone racing, which incorporates additional thrust signals to better capture drone motion.
Evaluation Metrics We use the following metrics for evaluation: 1) Absolute Translation Error (ATE, $\mathrm{m}$): the average Root Mean Squared Error (RMSE) between the estimated and ground-truth positions over all time points:
2) Relative Translation Error (RTE, $\mathrm{m}$): the average RMSE of the relative displacements over predefined time intervals.
Our evaluation adopts a 5-second interval configuration.
We define the improvement percentage as the relative reduction in error compared to the baseline. Specifically, it is computed as the difference between the error of baseline and the proposed method, divided by the baseline error.
Training Details We use the Adam optimizer with an initial learning rate of 0.001. A learning rate scheduler is implemented following the ReduceLROnPlateau, with a patience of 5 epochs and a decay factor of 0.2. The batch size is 128. Our training and testing window size is 1000 frames (equal to 5 seconds in the EuRoC dataset). In the CNN encoder, we include a dropout layer with $p=0.5$ to reduce overfitting.
IV-B Results
IV-B1 Evaluation on Blackbird Dataset
We employ the same sequences from Blackbird as used in IMO [6] and adopt their same train-test split configuration. We name these sequences as SEEN sequences because the training and testing data are from the same sequence. To evaluate generalization, we select five additional trajectories from BlackBird dataset as UNSEEN sequences, which are excluded from the training set. More details on separation are in Appendix. A-B1.